Hello,
I just began 9th grade, and tonight I received an assignment on quite a few of some word problems where I need to "Solve using the five-step plan. Write out each step. A choice of possible numbers for one unknown is given."
I've had much frustration and trouble with this... Here's *one of* the questions that I'm having trouble with
"The height of a flagpole is three fourths the height of the school. The difference in their heights is 4.5 m. What is the height of my school? Choices for school's height: 18,20,24."
My teacher wants me to do the four-step box way of showing and solving the equation. Such as:
1... Let s= School's height
Let (3/4)s= Flagpole's height
2... s-4.5 = 3/4s
3... (Solving the equation, showing work)
4... *solution*= School's height
*other solution*= Flagpole's Height
Thanks. I really need to complete this homework tonight, would appreciate your help please explain to me how to do this!
Thank you!
: Hello,
I just began 9th grade, and tonight I received an assignment on quite a few of some word problems where I need to "Solve using the five-step plan. Write out each step. A choice of possible numbers for one unknown is given."
I've had much frustration and trouble with this... Here's *one of* the questions that I'm having trouble with
"The height of a flagpole is three fourths the height of the school. The difference in their heights is 4.5 m. What is the height of my school? Choices for school's height: 18,20,24."
My teacher wants me to do the four-step box way of showing and solving the equation. Such as:
1... Let s= School's height
Let (3/4)s= Flagpole's height
2... s-4.5 = 3/4s
3... (Solving the equation, showing work)
4... *solution*= School's height
*other solution*= Flagpole's Height
Thanks. I really need to complete this homework tonight, would appreciate your help please explain to me how to do this!
Thank you!

Answer by Earlsdon(3719)

Well, your initial step is ok, but when you wrote step 2, you should have used parentheses, thus:
2...s-4.5 = (3/4)s
s-4.5 = (3/4)s So now you need to simplify this and solve for s, the height of your school.
s-4.5 = (3/4)s Add 4.5 to both sides of the equation.
s = 4.5+(3/4)s Now subtract (3/4)s from both sides.
s-(3/4)s = 4.5
(1/4)s = 4.5 Multiply both sides by 4 to clear the fraction.
s = 4*(4.5)
s = 18
So the height of your school is 18 meters.